• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jixie, Bob, Thai

    Log entry time 16:20:25 on April20,2012

    Entry number 376635

    keyword=Day shift summary

    8:00 Shift started. MCC is still tuning the beam.
    8:08 MCC said they were cycling the FZ magnet. Will delieve tune beam when finished.
    8:30 Halog was not posting any log. Adaql# machine did not accept ssh due to losing connection to home disk.
    Bob tried to fix them. IT expert Paul Letta came in to reboot the file server. All adaql# were rebooted by 9:30 then DAQ were tested.
    9:30 Javier told that the septa regulation was availabe. He shown how to bring the septa regulation screen.
    9:37 Beam is available. Start to determine the beam center.
    11:00 After one and a half hours of testing, We get to the conclusion that the beam is about 6mm below the horizontal at the target insert.
    We move the target insert 6 mm below.
    11:10 MCC took the beam away to change the wien angle. The wien angle was flipped to the wrong direction. It was not good for Hall A.
    12:20 MCC told they were setting up high current for Hall C. Beam will not be availabe in at least an hour. Move target to home.
    13:44 Change Right HRS to delta=0%. In order to save the time of draining LHe, I decided to do beam position scan before delta scan. Dustin will drain the LHe during take data for beam position scan.
    14:00 MCC told that A3 and A4 current to voltage converters for the injector were not working well. Need another hour to work on them.
    16:00 Shift ended, Still waiting for beam.